OpenStudy (anonymous):

There are 8 runners in a race. How many ways can the runners finish first, second, and third? A.8 B.24 C.336 D.6720

OpenStudy (anonymous):

you know permutations? 8*7*6=?

OpenStudy (anonymous):

Imagine you are GOD and first you will select one runner out of 8 for the first position then out of remaining 7 you will select the second and third from the remaining 6 So Ans = 8*7*6 = 336

OpenStudy (anonymous):

it can be solve without imagin she/he a God -.-

OpenStudy (anonymous):

Lol! Thank you both. @aviz great way of putting it!

OpenStudy (anonymous):

c

OpenStudy (anonymous):

=.=